can you help me find the inverse of the function g(x)= 1/(2x)
Answer by stanbon(57967)

inverse of the function g(x)= 1/(2x)
Interchange x and y to get:
x=1/(2y)
Solve for y as follows:
2y=1/(x)
y=1/(2x)
The function is its own inverse.
Cheers,
Stan H.